Grade 11Trigonometry

If value of Sin x +cos x =√2 then find sin x -cos x is equal to

Tanish
7 Years agoGrade 11

1 Answer

Aditya Gupta
7 Years ago
given that Sin x +cos x =√2
let Sin x – cos x = k
square and add both eqns
sin^2x+cos^2x+2sinxcosx + sin^2x+cos^2x – 2sinxcosx= 2+k^2
2= 2+k^2
or k=0
kindly approve :)